OpenStudy (anonymous):

Need Help PLEASE!!!!!!!!!! Write the equation of the line that is parallel to the line 3x − y = −3 and passes through the point (4, −2).

OpenStudy (anonymous):

NEED HELP PLEASE!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

3x - y = -3 (original) -3x + y = 3 (multiply the whole thing by -1 y = 3x + 3 (add or subtract to get y alone) your slope is 3, so m=3 you can substitute your point into (y - y1) = m (x - x1) to generate an equation for it. y + 2 = 3 ( x - 4 ) y + 2 = 3x - 12 y = 3x - 14

OpenStudy (anonymous):

@Zilther THANKS! THERE ARE MORE PROBLEMS LIKE THIS DO U MIND IF I TRY THEM ON MY OWN AND YOU COULD HELP ME CORRECT THEM AND TELL ME WHERE I WENT WRONG?

OpenStudy (anonymous):

if you say @zilther it will alert me and I can help you through more of them if you like. :)

OpenStudy (anonymous):

@@zilther THE QUESTIOIN IS : Write the equation of the line that is parallel to the line y = −3x + 12 and passes through the point (−1, 6)

OpenStudy (anonymous):

I found the slop m=-9 but im now stuck

OpenStudy (anonymous):

^@zilther

OpenStudy (anonymous):

@zilther

OpenStudy (anonymous):

okay, look at the formula that's up there ^ ( y - y1 ) = m ( x - x1 ) substitute your x and y vales for y1 and x1, then simplify and tell me what you get.

OpenStudy (anonymous):

your x1 and y1 values are the ones that you said the line passes through, ( -1 , 6 )

OpenStudy (anonymous):

parallel lines only share the same slope, the y-intercept doesn't matter at all, so keep that in mind.

OpenStudy (anonymous):

(y2-6) = -9 (x2-(-1) @zilther

OpenStudy (anonymous):

there's no y2, lol you end up with y - 6 = -9 ( x - (-1)) or y - 6 = -9 ( x + 1 ) if you simplify it. just follow PEMDAS and balance the equation. :)

OpenStudy (anonymous):

y = -6 = 9? @Zilther

OpenStudy (anonymous):

sorry for the late response. y - 6 = -9 ( x - (-1) ) y - 6 = -9 ( x + 1 ) <- simplified a little to get rid of the parenthesis you follow the distributive property -9 * x = -9x -9 * 1 = -9 so you end up with y - 6 = -9x - 9 after getting rid of the parenthesis y - 6 = -9x - 9 add 6 to each side to get rid of the -6 on the side of y y = -9x - 3 there's your equation.

OpenStudy (anonymous):

thats none of the answers tho Ay = one thirdx + 7 B. y = −3x + 3 C.y = one thirdx + 3 D. y = −3x + 7 @Zilther

OpenStudy (anonymous):

I have no idea what happened in that math. your slope is -3, and the point is ( -1 , 6 ) ( y - 6 ) = -3 ( x - (-1) ) y - 6 = -3 ( x + 1 ) y - 6 = -3x - 3 y = -3x + 3 <- that's the right answer

OpenStudy (anonymous):

sorry about that O.o;